A metallic wire of density d floats in water. They max. radius of the wire so that the wire may not sink will be(surface tension of water = T and angle of contact Q = o degree).
please explain.

Grade:12th pass

1 Answers

Vikas TU
7 years ago
We know the SURFACE TENSION as:
T = force/length 
or  = F/l
F =>  mg = g*PI*r^2*d*length 
calculating r value from here we get, 
r = root( m / g*pi*d*length)